Operation Summary for September 23, 2009 - September 25, 2009
Wednesday Sept. 23
The day shift began with Operations monitoring the Tevatron (TeV) store 7202 (49.26E30) and the Recycler stash (177.4E10), with the Antiproton Source stacking (16.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam even though MiniBooNE still had air conditioning problems, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
Operations terminated store 7202 at 9:10 AM.
The Neutron Therapy Facility (NTF) completed treatment at 12:13 PM on their last patient for day.
Operations prepared Linac for access at 12:23 PM.
Operations prepared Booster for access at 1:34 PM.
Operations reported at 8:36 PM that vacuum technicians had found and fixed a leak on a small bit of corroded stainless steel beam pipe in MI-30.
MI, Pbar, MiniBooNE, and NuMI resumed taking beam by 9:45 PM.
At 11:33 PM, Operation reported having problems with the MI-52 kicker, which was causing the P1 transfer line beam permit to trip off. Operators contacted kicker personnel. This problem halted stacking.
Thursday Sept. 24
The midnight shift began with Operations monitoring store 7203 (49.83E30) and the stack (48.1E10), with no stash or stacking due to kicker (KPS52) problems, and with NuMI and MiniBooNE taking beam.
At 12:25 AM, a knowledgeable kicker person recommended operators try to dry out the condensation on the kicker racks.
At 1:28 AM, operators reported drying the MI kicker power supply (KPS52) with a heat gun. Operators asked the Duty Mechanic to lower the thermostat for the kicker room to make the AC come on.
At 1:32 AM, Operations reported noise problems with two Linac quadrupole power supplies (L:QTM3C and QTM3CS). Operators put a fan on the power supplies.
Operations reported at 2:33 AM that they had stashed antiprotons into the Recycler.
A MI kicker person reported at 2:42 AM that he had positioned two heaters and a warm air blower on the kicker power supply about an hour earlier and things were drying out.
At 3:26 AM, Operations reported that a power cycle of the Linac quadrupole power supplies resolved the noise problems.
Operations terminated store 7203 at 7:03 AM.
Operations prepared the TeV, CDF, and D0 for access at 7:16 AM.
CDF and D0 were out of access by 10:56 AM.
The TeV came out of access and Operations began the TeV turn-on at 12:26 PM.
Operation put the TeV through a dry squeeze at 3:27 PM.
NTF finished treating their last patient for the day at 4:01 PM.
Operations put the TeV into shot setup at 5:11 PM and established store 7206 at 6:34 PM with an AIL of 179.93E30.
The TeV RF person brought the APS back online. He could not find anything wrong with the station or supply.
Friday Sept. 25
The midnight shift began with Operations monitoring the stash (141.3E10), with Pbar stacking (25.6E10), and with NuMI and MiniBooNE taking beam.
Operations put the TeV through a dry squeeze at 12:42 AM.
Operations put the TeV into shot setup and established store 7209 at 2:19 AM with an AIL of 120.6E30.
The Plans for Friday and the WeekendThe plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 09/25/2009 email Fermilab|
|Security, Privacy, Legal|